Two friends live 7 miles apart. One Saturday, the two friends set out on their bikes at 8 am and started riding towards each other. One rides at 0.2 miles per minute, and the other rides at 0.15 miles per minute. At what time will the two friends meet?
Answer by ikleyn(52788)

.

The approaching rate is 0.15 + 0.2 = 0.35 miles per minute.


Hence, they will meet in 7/0.35 = 700/35 = 20 minutes.


ANSWER.  They will meet at 8:20 am. 

Solved.